OpenStudy (anonymous):

what is 0.08333333333 as a fraction?

OpenStudy (ksaimouli):

8333333333/10^11

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

don you mean 0.08333333333...?

OpenStudy (anonymous):

if so then 0.08333... = 0.08 + 0.00333... = 0.08 + 3/900 = 1/12

OpenStudy (anonymous):

yes it's 1/12